OpenStudy (anonymous):

At a certain time of day, the angle of elevation of the sun is 30°. A tree has a shadow that is 25 feet long. Find the height of the tree to the nearest foot. A) 13 ft B) 14 ft C) 22 ft D) 29 ft

OpenStudy (anonymous):

the shadow goes along the ground, and the tree is straight up. So we have angle =40, adjacent side = 35, opposite side = x. tan(40) = x/35 .8391 = x/35 x = 29.37 So your answer is D.(: